# A novel sensitivity analysis method for agent-based models stratifies in-silico tumor spheroid simulations

### v2

@@ Abstract
Agent-based models (ABMs) are widely used in biology to understand how individual actions scale into emergent population behavior. Modelers employ sensitivity analysis (SA) algorithms to quantify input parameters’ impact on model outputs, however, it is hard to perform SA for ABMs due to their computational and complex nature. In this work, we develop the Simulate, Summarize, Reduce, Cluster, and Analyze (SSRCA) methodology, a machine-learning based pipeline designed to facilitate SA for ABMs. In particular, SSRCA can achieve the following tasks for ABMS: 1) identify sensitive model parameters, 2) reveal common output model patterns, and 3) determine which input parameter values generate these patterns. We use an example ABM of tumor spheroid growth to showcase how SSRCA provides similar SA results to the popular Sobol’ Method while also identifying four common patterns from the ABM and the parameter regions that generate these outputs. This analysis could streamline data-driven tasks, such as parameter estimation, for ABMs by reducing parameter space. While we highlight these results with an ABM on tumor spheroid formation, the SSRCA methodology is broadly applicable to biological ABMs.
@@ 1 Introduction
Multicellular tumor spheroids are a widely-used experimental procedure to study avascular tumor growth and anti-cancer drug therapies [1, 2, 3]. As opposed to 2-dimensional assay experiments, 3-dimensional tumor spheroid experiments closely mimic the in vivo tumor microenvironment with realistic cell-cell interactions, chemical gradients, the formation of hypoxic regions, etc. These conditions lead to high levels of intratumoral heterogeneity. For example, cells in oxygen rich environments will rapidly divide and increase in number, whereas those in hypoxic regions may become quiescent or die. The fluorescent ubiquitination-based cell cycle indicator (FUCCI) imaging technique is a useful modality to measure and quantify tumor spheroid heterogeneity by tracking individual cells’ progression through the cell cycle in real-time [3, 4, 5]. Using FUCCI, the nuclei of cells in G1 phase fluoresce red, while the nuclei of cells in G2/S/M phase fluoresce green. The nuclei of cells in early S (eS) phase fluoresce yellow due to the presence of both red and green fluorescence [4].
@@
Mathematical models provide an efficient and cost-effective approach to generate hypotheses for tumor spheroid formation and optimally design future experiments [6, 7]. Common approaches for simulating in-silico tumor spheroids include continuum models [8, 9, 10, 11, 12, 13] and cellular automata [14, 15, 16]. Agent-based models (ABMs) are another popular modeling choice due to their ability to capture the discrete and stochastic nature of tumor spheroids [17, 18, 19, 14, 20]. In an ABM, modelers simulate pre-defined individual-level rules to determine how they translate into collective behavior. For example, Klowss et al. developed a hybrid ABM that consists of both discrete tumor cells and a continuous variable representing a nutrient (such as oxygen) that controls cells’ progression through the cell cycle [21]. This model includes rules dictating cell migration, death, and transition between cell cycle states. The resulting in silico tumor spheroid captures many realistic facets of in vitro tumor spheroid growth: an outer proliferating shell, an inner growth-arrested shell, and a necrotic core [22]. Model simulations serve to precisely predict how experimental conditions and/or design choices (including initial spheroid size, kinetic parameter rates, time duration, etc.) impact these patterns. In another study, Rocha et al. demonstrated that their epithelial metastasis ABM could be used as a cancer patient digital twin to develop patient-specific models for precision medicine treatment plans [23].
@@
The analysis of ABMs presents many challenges; notably, ABMs often require many parameter values, are not amenable to extensive simulation due to long simulation times, and their outputs are complex and noisy. These difficulties hinder the application of standard modeling techniques for ABMs, including parameter estimation, uncertainty quantification, and sensitivity analysis. As one example, consider the 3-dimensional Klowss Model from [21]: There are 25 model parameters: 13 can be estimated from biological experiments, 2 can be determined from numerical experiments, leaving 10 unknown parameters. Impressively, the authors selected specific values for all 10 parameters and demonstrated excellent agreement between their in-silico simulations and in-vitro spheroid experiments using human primary melanoma cells. However, there is no guarantee that the same values would generalize to data from other cell lines. Estimating all 10 parameter values from new data would require exhaustively searching the 10-dimensional parameter space, which is infeasible due to long simulation times and the curse of dimensionality [24]. In addition to parameter estimation, these obstacles inhibit our ability to perform feature mapping, which links the model’s input parameter values to its output patterns. In the context of tumor spheroids, these limitations prevent the use of complex models for precision medicine diagnostics [25].
@@
Parameter sensitivity analysis (SA) is an area of research to identify a model’s most influential parameters. Many aspects of the modeling process are encompassed in SA, including quantifying the uncertainty of model outputs, determining influential input parameters, and feature mapping [26, 27]. Modelers can fix the insensitive parameters to reduce the dimensionality of the input parameter space while still capturing the main model behavior. There are many different approaches to performing SA [28, 29]. For example, the Morris one-at-a-time (MOAT) method is a computationally-efficient algorithm that varies one parameter at a time. This simple approach ignores interactions between parameters and can thus generate misleading results for complex models. The Sobol’ Method is a global SA (GSA) approach that ranks parameters based on their contribution to the model’s variance using computed sensitivity index values. These Sobol’ index values incorporate parameter interactions, but this added information comes at the cost of higher computational expenses. Regression-based methods, such as partial rank correlation coefficient, are another popular SA method in which one fits a regression model to predict the model’s output values from its input parameters. These approaches can be used to perform feature mapping, however, it can be challenging to determine a suitable regression model for nonlinear models [30].
@@
Despite the wide and successful use of SA methods, their application to ABMs has been limited. For example, feature mapping is challenging to perform for an ABM because nearby parameter values can yield divergent patterns. Broeke et al. assessed the performance of MOAT and the Sobol’ Method for ABMs, and found that the simple MOAT approach can be used for preliminary feature mapping analysis, but that the Sobol’ Method cannot because its analyses are computed over the entire parameter space instead of considering distinct parameter regions [26]. Other challenges for performing SA for ABMs arise due to their time-varying, nonlinear, computational, and multi-scale natures [31, 27, 30]. Instead of applying standard GSA methods to ABMs, there is an urgent need for the development of novel ABM-specific SA approaches to fully harness the potential of ABMs for modeling complex phenomena.
Towards this end, we introduce the Simulate, Summarize, Reduce, Cluster, and Analyze (SSRCA, pronounced “circa”) methodology to perform GSA for ABMs. An advantage of this approach over other GSA algorithms, such as the Sobol’ indices method, is that it can both identify sensitive parameters and perform feature mapping. In particular, SSRCA can achieve the following tasks: 1) identifies sensitive model parameters, 2) reveals common output model patterns, and 3) determines which parameter values generate these patterns. We focus our work on a 2-dimensional verision of the Klowss Model from [21] using two simulated datasets. For a small dataset (generated by varying 2 model parameters), SSRCA directly links parameter values to one of four typical model patterns. For a larger dataset (generated by varying 10 model parameters), SSRCA classifies four sensitive model parameters, and estimates the marginal parameter distributions underlying each such pattern. These results suggest that SSRCA can describe the parameter values (and, in turn, biological mechanisms) underlying each output model pattern for complex ABMs. This study is focused on a model of tumor spheroid growth, however, the SSRCA methodology is broadly applicable to biological ABMs over all scales, ranging from intracellular protein dynamics to disease spread in human populations [32, 33].
@@ 2 Methods
In this section, we review the Klowss model from [21] 2.1, the SSRCA methdology 2.2, and the Sobol’ Method 2.3.
@@ 2.1 The Klowss Model of tumor spheroid growth
We implement our proposed methodology on the Klowss Model, an off-lattice hybrid ABM of tumor spheroid growth [21]. We provide details on the model and its implementation in Appendix A. Briefly, the model contains $N(t)$ agents that are colored according to their cell cycle status over time: agents in the G1 phase are colored red, agents in the early S stage are yellow, agents in the S/G2/M phase are green, and dead cells are blue (Figure 1(a)). All agents perform rules on migration, death, and cell cycle progression; when an agent transitions from green to red, it creates an additional red daughter cell to imitate cell division. (Figure 1(b-d)). The kinetic rates for each rule depend on a continuous variable, $c(\bm{x},t)$, such as oxygen (Figure 1(e)). The spatiotemporal concentration of $c(\bm{x},t)$ varies due to diffusion and agent consumption.
Figure 1: Tumor spheroid model overview. (a) Agent cells are colored according to their cell cycle status. (b) All agents migrate with rate $m(c)$, and move the length of a cell diameter in a randomly chosen direction. (c) All agents die with rate $d(c)$ and change their cell cycle status to dead. (d) The cell cycle is modeled as follows: red agents transition to yellow with rate $R_{r}(c)$, yellow agents transition to green with rate $R_{y}(c)$, green agents undergo mitosis with rate $R_{g}(x)$ in which they become red and create a daughter red cell; the daughter cell is placed one cell diameter away in a randomly chosen direction. (e) All model rates depend on the nutrient available at the agent’s physical location.
The Klowss Model from [21] is 3-dimensional in space, however, we find that simulating the model in 2-dimensions can provide good agreement with the middle cross-section of a 3-dimensional simulation. As a computationally-efficient alternative to this 3-dimensional model, we compute the model in 2-dimensions ($\bm{x}_{n}(t)=(x_{n}(t),y_{n}(t))^{T}$) to ease the model’s computational requirements. This is a common simplification used when performing large-scale ABM simulations [23, 34, 35, 36].
2.2 The SSRCA methodology
  1. 1.
    Simulate model data (Section 2.2.1),
  2. 2.
    Summarize simulations (Section 2.2.2),
  3. 3.
    Reducing dimension (Section 2.2.3),
  4. 4.
    Cluster data (Section 2.2.4), and
  5. 5.
    statistical Analysis (Section 2.2.5).
2.2.1 Simulate model data
We simulate the model over a sample of parameter values from a specified distribution. In this study, we apply the SSRCA methodology to two simulated datasets to illustrate its performance on multiple data scales. A brief description of each parameter, including which parameters are varied and the minimum and maximum values of all varied parameters, is provided in Table 1. For simplicity, all parameters were varied between half of their baseline value and twice their baseline value.
@@ 2.2.1 Simulate model data
Table 1: ABM parameters. For all the perturbed parameters, the minimum parameter value is always half the baseline value, while the maximum parameter value is always double the baseline value.
Dataset 1: In this dataset, we vary the hill function index for arrest, $\eta_{1}$, and the critical arrest concentration, $c_{a}$, and set all other parameters to their base values from Table 1. We vary both $\eta_{1}$ and $c_{a}$ over 11 logarithmically spaced values that range over their minimum and maximum values, totaling 121 parameter combinations. We perform 10 model simulations for each combination, resulting in $11\times 11\times 10=1,210$ total model simulations.
Dataset 2: In this dataset, we vary two parameters at a time while fixing all others at their base values. We consider all (${10\choose 2}=45$ pairs of the varying parameters from Table 1. For each pair, we vary both parameters over 11 logarithmically spaced values that range over their minimum and maximum values, totaling 121 parameter combinations for each pair of parameters. We perform 10 model simulations for each parameter combination, leading to $45\times 11\times 11\times 10=54,450$ total simulations.
@@ 2.2.2 Summarize simulations
Recording the locations and cell cycle states of all $N(t)$ agents over time results in prohibitively large datasets because each simulation contains thousands of agents. Furthermore, standard machine learning algorithms require fixed length vectors, but each simulation has a different number of cell counts. For these reasons, we compute fixed length descriptor vectors (DVs) to concisely summarize each model simulation. We consider two separate DVs in this study: cell subpopulation counts over time, and the final living cell density over space. Other DVs are possible; the appropriate DV choice depends on the model approach and application [34, 36].
[…]
Final Cell Density: We create a two-dimensional histogram of living agent (those colored red, yellow, and green) locations at the final timepoint ($t=240$ hours) using binning values $\bm{x}^{\text{bin}}=\bm{y}^{\text{bin}}=\{0,40,80,\dots,1000\mu\text{m}\}$. We divide each histogram value by the binned area to estimate the spatial cell density. We vectorize each 2-dimensional density of size $25\times 25$ into a final vector of length 625.

### v3

# SSRCA: a novel machine learning pipeline to perform sensitivity analysis for agent-based models

@@ Abstract
Agent-based models (ABMs) are widely used in biology to understand how individual actions scale into emergent population behavior. Modelers employ sensitivity analysis (SA) algorithms to quantify input parameters’ impact on model outputs, however, it is hard to perform SA for ABMs due to their computational and complex nature. In this work, we develop the Simulate, Summarize, Reduce, Cluster, and Analyze (SSRCA) methodology, a machine-learning based pipeline designed to facilitate SA for ABMs. In particular, SSRCA can achieve the following tasks for ABMS: 1) identify sensitive model parameters, 2) reveal common output model patterns, and 3) determine which input parameter values generate these patterns. We use an example ABM of tumor spheroid growth to showcase how SSRCA identifies four common patterns from the ABM and the parameter regions that generate these outputs. Additionally, we compare the SA results between SSRCA and the popular Sobol’ Method and find that SSRCA’s identified sensitive parameters are robust to the choice of model descriptors while Sobol’s are not. This analysis could streamline data-driven tasks, such as parameter estimation, for ABMs by reducing parameter space. While we highlight these results with an ABM on tumor spheroid formation, the SSRCA Methodology is broadly applicable to biological ABMs.
@@ 1 Introduction
Mathematical models provide an efficient and cost-effective approach to generate hypotheses for biological processes and optimally design experiments [byrne_dissecting_2010, wallace_properties_2013]. Common approaches for simulating in-silico experiments include continuum models [greenspan_models_1972, MCELWAIN1977267, ward_mathematical_1997, 10.1093/imammb/18.2.131, Grimes, murphy_growth_2023] and cellular automata [MALLET2006334, bruningk_cellular_2019, messina_hybrid_2023]. Agent-based models (ABMs) are another popular modeling choice due to their ability to capture the discrete and stochastic nature of biological processes [10.1371/journal.pcbi.1006469, 10.1371/journal.pcbi.1007961, Cooper2020, MALLET2006334, rigid_body]. In an ABM, modelers simulate pre-defined individual-level rules to determine how they translate into collective behavior. For example, Klowss et al. developed a hybrid ABM that consists of both discrete tumor cells and a continuous variable representing a nutrient to imitate tumor spheroid experiments [4D_Tumor_Model]. This model includes rules dictating cell migration, death, and transition between cell cycle states. Model simulations serve to precisely predict how experimental conditions and/or design choices (including initial spheroid size, kinetic parameter rates, experimental duration, etc.) impact the final model outcomes. In another study, Rocha et al. demonstrated that their ABM could be used as a cancer patient digital twin to develop patient-specific models for precision medicine treatment plans [l_rocha_multiscale_2024].
@@
Despite the wide use of ABMs, their analysis presents many challenges; ABMs typically require many parameter values, are not amenable to extensive simulation due to long simulation times, and their outputs are complex and noisy. These difficulties hinder the application of standard modeling techniques for ABMs, including parameter estimation, uncertainty quantification, and sensitivity analysis. As one example, consider the Klowss Model from [4D_Tumor_Model]: There are 25 model parameters: 13 can be estimated from biological experiments, 2 can be determined from numerical experiments, leaving 10 unknown parameters. Impressively, the authors in this study found values for all 10 parameters and demonstrated excellent agreement between their in-silico simulations and in-vitro spheroid experiments using human primary melanoma cells. However, there is no guarantee that the same values would generalize to data from other cell lines. Estimating all 10 parameter values from new data would require exhaustively searching the 10-dimensional parameter space, which is infeasible due to long simulation times and the curse of dimensionality [fernandez-martinez_curse_2020]. In addition to parameter estimation, these obstacles inhibit our ability to perform feature mapping, which links the model’s input parameter values to its output patterns.
@@
Parameter sensitivity analysis (SA) is an area of research to identify a model’s most influential parameters. Many aspects of the modeling process are encompassed in SA, including quantifying the uncertainty of model outputs, determining influential input parameters, and feature mapping [ten_broeke_which_2016, bergman_efficient_2025]. Modelers can fix the insensitive parameters to reduce the dimensionality of the input parameter space while still capturing the main model behavior. There are many different approaches to performing SA [dellino_review_2015, pianosi_sensitivity_2016]. For example, the Morris one-at-a-time (MOAT) method is a computationally-efficient algorithm that varies one parameter at a time. This simple approach ignores interactions between parameters, however, and can thus generate misleading results for complex models. The Sobol’ Method is a global SA (GSA) approach that ranks parameters based on their contribution to the model’s variance using computed sensitivity index values. These Sobol’ index values incorporate parameter interactions, but this added information comes at the cost of higher computational expenses. Regression-based methods, such as partial rank correlation coefficient, are another popular SA method in which one fits a regression model to predict the model’s output values from its input parameters. While regression-based approaches can be used to perform feature mapping, it can be challenging to determine a suitable regression model for nonlinear models [ligmann-zielinska_one_2020].
@@
Despite the wide and successful use of SA methods, their application to ABMs has been limited. For example, feature mapping is challenging to perform for an ABM because nearby parameter values can yield divergent patterns. Broeke et al. assessed the performance of MOAT and the Sobol’ Method for ABMs, and found that the simple MOAT approach can be used for preliminary feature mapping analysis, but that the Sobol’ Method cannot because its analyses are computed over the entire parameter space instead of considering distinct parameter regions [ten_broeke_which_2016]. Other challenges for performing SA for ABMs arise due to their time-varying, nonlinear, computational, and multi-scale natures [bergman_modeling_2021, bergman_efficient_2025, ligmann-zielinska_one_2020]. Instead of applying standard GSA methods to ABMs, there is an urgent need for the development of novel ABM-specific SA approaches to fully harness the potential of ABMs for modeling complex phenomena. Towards this end, we introduce the Simulate, Summarize, Reduce, Cluster, and Analyze (SSRCA, pronounced “circa”) methodology to perform GSA for ABMs (Figure 1). An advantage of this approach over other GSA algorithms, such as the Sobol’ indices method, is that it can both identify sensitive parameters and perform feature mapping. In particular, SSRCA can achieve the following tasks: 1) identifies sensitive model parameters, 2) reveals common output model patterns, and 3) determines which parameter values generate these patterns.
We focus our application of the SSRCA Methodology on the Klowss Model from [4D_Tumor_Model]. This 3-dimensional model imitates tumor spheroid experiments, which contain high levels of intratumoral heterogeneity. For example, cells in oxygen rich areas will rapidly divide and increase in number, whereas those in hypoxic regions may become quiescent or die. The fluorescent ubiquitination-based cell cycle indicator (FUCCI) imaging technique allows experimentalists to measure and quantify population heterogeneity by tracking individual cells’ progression through the cell cycle in real-time [koledova_real-time_2017, FUCCI, Vittadello]. Simulations of the Klowss Model capture many realistic facets of in vitro tumor spheroid growth: an outer proliferating shell, an inner growth-arrested shell, and a necrotic core [murphy_designing_2022]. In this study, we apply the SSRCA Methodology to a 2-dimensional version of the Klowss Model from [4D_Tumor_Model] using two simulated datasets. For a small dataset (generated by varying 2 model parameters), SSRCA directly links parameter values to one of four typical model patterns quantifying the sizes of the proliferating shell and necrotic core. For a larger dataset (generated by varying 10 model parameters), SSRCA identifies four sensitive model parameters, each of which is involved in cell cycle entry or death, and the parameter distributions underlying the common model patterns. These results suggest that cell cycle entry and death are the most impactful processes involved in tumor spheroid formation, and SSRCA reveals the possible model outputs when varying these sensitive parameters. We highlight the SSRCA Methodology’s capability for performing GSA on a tumor spheroid ABM in this work, however, this approach is broadly relevant to biological ABMs with applications in intracellular protein dynamics, disease spread, and ecology [ciocanel_simulated_2022, kim_simulation_2022, bernoff_agent-based_2020].
@@ 2 Methods
In this section, we review the Klowss Model from [4D_Tumor_Model] in Section 2.1, detail the SSRCA Methdology in Section 2.2, describe how we re-parameterize the Klowss Model in 2 spatial dimensions in Section 2.3, and review the Sobol’ Method in Section 2.4.
@@ 2.1 The Klowss Model of tumor spheroid growth
Many mathematical models of tumor spheroid dynamics have been developed previously [karolak_towards_2018], with most capturing key characteristics of in vitro tumor spheroid growth: the formation of a necrotic core in the spheroid’s center and an outer proliferating ring due to nutrient diffusion and consumption [murphy_designing_2022]. Within these modeling studies, many different modeling approaches have been implemented to describe tumor cell locations or densities over time, including cellular automata [MALLET2006334, messina_hybrid_2023, 10.1371/journal.pcbi.1006469, bruningk_cellular_2019, song_chen_agent-based_2004], off-lattice ABMs [10.1371/journal.pcbi.1007961, 4D_Tumor_Model, milotti_emergent_2010], Voronoi-latticed ABMs [cleri_agent-based_2019, jagiella_inferring_2016], or differential equation models [greenspan_models_1972, murphy_designing_2022, murphy_growth_2023]. In this study, we choose to focus our application of the SSRCA methodology on the off-lattice Klowss Model from [4D_Tumor_Model] for two key reasons. First, this model has already been calibrated to in vitro experimental data previously, which indicates it captures realistic biological behavior. Second, this model is of moderate complexity with 25 parameters (10 of which were assumed in [4D_Tumor_Model] without biological justification). This provides us with an opportunity to determine which of these 10 assumed parameters most significantly impact the model’s outputs and to explore how the model’s dynamics change when varying the model’s sensitive parameters. Even though a 3-dimensional model was presented in [4D_Tumor_Model], we simulate the model in 2 spatial dimensions to ease the computational requirements; see Section 2.3 for more details.
We provide details on the Klowss Model and our implementation in Appendix S1 and the baseline parameter values used in Table 1. Briefly, the model contains $N(t)$ agents. Each agent may be in one of three living states: G1, early S, or G2/S/M (Figure 2(a)). All living agents perform rules on migration, death, and cell cycle progression; agents undergo mitosis by creating a daughter cell in the G1 state when transitioning from the S/G2/M state to the G1 state (Figure 2(a)). Additionally, cells can transition from any of the three living states into the dead state (Figure 2(a)). The kinetic rates for each rule depend on a continuous variable, $c(\bm{x},t)$, such as oxygen (Figure 2(b)). The rates of cell migration, $m(c)$, and entry into the cell cycle, $R_{r}(c),$ are both modeled using Hill functions that increase with $c(\bm{x},t)$; the death rate, $d(c)$, is modeled using a Hill function that decreases with $c(\bm{x},t)$; and the rates from early S to S/G2M, $R_{y}(c)$, and for mitosis, $R_{g}(c)$, are both constant. The spatiotemporal concentration of $c(\bm{x},t)$ varies due to diffusion and agent consumption, see Appendix S1.2. We describe the initialization process for the model in Appendix S1.3
@@
Table 1: ABM parameters. For all the perturbed parameters, the minimum parameter value is always half the baseline value, while the maximum parameter value is always double the baseline value. All baseline values are from [4D_Tumor_Model], except $N(0)$ and $\alpha$, which are determined in Section 2.3. For more detailed parameter information, see Section S1.
Figure 2: Klowss model overview and associated descriptor vectors. (a) The cell cycle is modeled as follows: agents in the G1 state transition to the early S state with rate $R_{r}(c)$, agents in the early S state transition to the S/G2/M state with rate $R_{y}(c)$, agents in the S/G2/M state undergo mitosis with rate $R_{g}(x)$ in which they become red and create a daughter red cells in the G1 state; the daughter cell is placed one cell diameter away in a randomly chosen direction. All agents die with rate $d(c)$ and change their cell cycle status to dead. Not shown: All agents migrate with rate $m(c)$, and move the length of a cell diameter in a randomly chosen direction. (b) All model rates depend on the nutrient available (in arbitrary units (AU)) at the agent’s physical location. The depicted rates qualitatively illustrate cells’ behavior in the model but are not the precise rates used for the simulations. (c) Snapshots at times t = 0, 2, 6, and 10 days from the baseline simulation. d) The Cell counts DV for the baseline simulation. e) The Final cell density DV for the baseline simulation. The solid (dashed) circle has a radius of 137 $\mu m$ (322 $\mu m$).
2.2 The SSRCA Methodology
  1. 1.
    Simulate model data (Section 2.2.1),
  2. 2.
    Summarize simulations (Section 2.2.2),
  3. 3.
    Reduce dimension (Section 2.2.3),
  4. 4.
    Cluster data (Section 2.2.4), and
  5. 5.
    statistical Analysis (Section 2.2.5).
2.2.1 Simulate model data
We simulate the model over a sample of parameter values from a specified distribution. In this study, we apply the SSRCA Methodology to two simulated datasets to illustrate its performance on both a small and large dataset. A brief description of each parameter, including which parameters are varied and the minimum and maximum values of all varied parameters, is provided in Table 1. For simplicity, all parameters were varied between half of their baseline value and twice their baseline value.
The simple 2-parameter dataset: In this dataset, we vary the Hill function index for arrest, $\eta_{1}$, and the critical arrest concentration, $c_{a}$, and set all other parameters to their base values from Table 1. We vary both $\eta_{1}$ and $c_{a}$ over 11 logarithmically spaced values that range over their minimum and maximum values, totaling 121 parameter combinations. We perform 10 model simulations for each combination, resulting in $11\times 11\times 10=1,210$ total model simulations.
The large 10-parameter dataset: In this dataset, we vary parameters two at a time while fixing all others at their base values. We consider all ${10\choose 2}=45$ pairs of the varying parameters from Table 1. For each pair, we vary both parameters over 11 logarithmically spaced values that range over their minimum and maximum values, totaling 121 parameter combinations for each pair of parameters. We perform 10 model simulations for each parameter combination, leading to $45\times 11\times 11\times 10=54,450$ total simulations.
We used The College of New Jersey’s Electronic Laboratory for Science and Analysis (ELSA) high performance computer. We computed the large 10-parameter dataset using 450 1-core SLURM jobs, each with 6GB of RAM. Each job ran on a node consisting of a dual Intel(R) Xeon(R) Gold 6130 CPU processer with 192 total GB of RAM. All 450 jobs completed within 24 hours.
@@ 2.2.2 Summarize simulations
Recording the locations and cell cycle states of all $N(t)$ agents over time results in prohibitively large datasets because each simulation contains thousands of agents. Furthermore, standard machine learning algorithms require fixed length vectors, but each simulation has a different number of cell counts. For these reasons, we compute fixed length descriptor vectors (DVs) to concisely summarize each model simulation. We consider two separate DVs in this study: cell subpopulation counts over time, and the final living cell density over space. Other DVs are possible; the appropriate DV choice depends on the model approach and application [angio_tda, bhaskar_analyzing_2019].
Cell counts: We count the number of red, yellow, green, and dead cells over time as $N_{r}(\bm{t})$, $N_{y}(\bm{t})$, $N_{g}(\bm{t})$, and $N_{d}(\bm{t})$, respectively, for the 241 timepoints at $\bm{t}=\{0,1/24,2/24,\dots,10\text{ days}\}$. For the final DV, we concatenate these four vectors together into a vector of length $4\times 241=964$.
Final Cell Density: We create a two-dimensional histogram of living agent (those colored red, yellow, and green) locations at the final timepoint ($t=10$ days) using binning values $\bm{x}^{\text{bin}}=\bm{y}^{\text{bin}}=\{0,40,80,\dots,1000\mu\text{m}\}$. We divide each histogram value by the binned area to estimate the spatial cell density. We vectorize each 2-dimensional density of size $25\times 25$ into a final vector of length 625.